# thanks: Corresponding Author (shared)thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: Corresponding Author (shared)

### v1

@@ IV Systems of $N\geq 2$ Complexified Oscillators
In how far do the phenomena uncovered for $N=2$ units transfer to larger systems? Here we first consider key aspects for $N\geq 2$ coupled complexified Kuramoto oscillators. Rewriting Eq. (3) as
@@
for each $n\in[N]$. For example, considering a system of identical complexified oscillators, i.e., $\omega_{n}=0$ for $n\in[N]$, the phase-lag parameter $\alpha$ delineates the boundary separating regions characterized by linear stability and instability of uniform oscillations, referred to as complete synchronization of complexified Kuramoto oscillators. Here, such instability point occurs at the argument $\alpha=\frac{\pi}{2}$ of the complex coupling $K$. Interestingly, the instability occurs at the same value $\alpha=\frac{\pi}{2}$ in a system of real-valued Kuramoto-Sakaguchi oscillatorssle .
@@
Figure  5 depicts time-averaged Kuramoto order parameters, directly measured from numerical simulations, as a function of the modulus $|K|$ of the coupling $K=|K|e^{\textrm{i}\alpha}$ in four different settings. First, for real coupling (with $\alpha=0$) and initial conditions with zero imaginary parts ($y_{n}(0)=0$ for all $n\in[N]$), a transition occurs as in the real-variable Kuramoto model. With increasing $|K|$, a partially locked state emerges from an incoherent state (red open diamonds). Indeed, the oscillators are constrained onto the real invariant manifold (5), so the dynamics exactly resembles that of the original, real-valued Kuramoto model, cf. also Fig. 1a. In contrast, the same system started from initial conditions with non-zero imaginary parts ($y_{n}(0)\neq 0$ for all $n\in[N]$) displays markedly different features (cf. Ref. Thümler et al., 2023): The Kuramoto order parameter discontinuously jumps from indicating an incoherent state coexisting with unstable complex-locked states to a stable complex-locked state, i.e., a stable fixed point in complex vector space defined in Ref. Thümler et al., 2023.
For generic complex coupling $\alpha=\frac{\pi}{2}-0.2$ (purple circles in Fig. 5) the system exhibits a discontinuous synchronization transition at substantially smaller $|K|$, where the majority of oscillators become closely locked and $r$ close to one. Past the transition, pronounced synchrony emerges with a distribution variance in the real parts $x_{n}$ on the order of $10^{-3}$ and non-zero imaginary parts. Finally, for a purely imaginary coupling with $\alpha=\frac{\pi}{2}$, the oscillators are incoherently drifting, yielding a low order parameter relative to other scenarios even at substantially large $|K|$ (green open squares). These findings qualitatively align with those for the two-oscillator systems discussed in Sec. III.3, where the systems fail to exhibit any classical locked state and are asynchronous for purely imaginary coupling, whereas oscillators subject to a generic complex coupling become locked at a stable fixed point and exhibit a substantial degree of synchrony.
@@ V Conclusions and Open Questions
More specifically, we have presented a range of results: For $N=2$ coupled units and purely imaginary coupling, we discovered motions of libration around fixed points of in-phase and antiphase states which are respectively local minima and maxima of a conserved quantity, and a ceaseless rotation along the cylindrical axis with an infinite winding number $q=\infty$. No attractive locked state emerges even for arbitrarily large $|K|$. In contrast, for generic complex coupling, trajectories rotate a few times, exhibit finite winding numbers $q<\infty$, and stop at a stabilized fixed point. We have analytically determined the winding number and derived its asymptotic scaling as $\alpha\rightarrow\pi/2$ from below, where the coupling becomes purely imaginary. For small $\beta=\pi/2-\alpha$, synchrony stabilizes: Slightly turning on the real part of a coupling, the neutrally stable fixed points of in-phase synchrony become stabilized to first order in $\beta$. Larger, finite-size systems of $N>2$ units similarly exhibit a classically locked state for purely imaginary coupling whereas a discontinuous transition emerges to a locked state for a generic complex coupling. The analyses above catalyze a range of future research.

### v2

@@ IV Systems of $N\geq 2$ Complexified Oscillators
How do the phenomena uncovered for $N=2$ units transfer to larger systems? Here we first consider key aspects for $N\geq 2$ coupled complexified Kuramoto oscillators. Rewriting Eq. (3) as
@@
for each $n\in[N]$. For example, considering a system of identical complexified oscillators, i.e., $\omega_{n}=0$ for $n\in[N]$, the phase-lag parameter $\alpha$ delineates the boundary separating regions characterized by linear stability and instability of uniform oscillations, referred to as complete synchronization of complexified Kuramoto oscillators. Here, such instability point occurs at the argument $\alpha=\frac{\pi}{2}$ of the complex coupling $K$. Interestingly, the instability occurs at the same value $\alpha=\frac{\pi}{2}$ in a system of real-valued Kuramoto-Sakaguchi oscillators.
@@
Figure  5 depicts time-averaged Kuramoto order parameters (2), directly measured from numerical simulations, as a function of the modulus $|K|$ of the coupling $K=|K|e^{\textrm{i}\alpha}$ in four different settings. First, for real coupling (with $\alpha=0$) and initial conditions with zero imaginary parts ($y_{n}(0)=0$ for all $n\in[N]$), a transition occurs as in the real-variable Kuramoto model. With increasing $|K|$, a partially locked state emerges from an incoherent state (red open diamonds). Indeed, the oscillators are constrained onto the real invariant manifold (5), so the dynamics exactly resembles that of the original, real-valued Kuramoto model, cf. also Fig. 1a. In contrast, the same system started from initial conditions with non-zero imaginary parts ($y_{n}(0)\neq 0$ for all $n\in[N]$) displays markedly different features (cf. Ref. Thümler et al., 2023): The Kuramoto order parameter discontinuously jumps from indicating an incoherent state coexisting with unstable complex-locked states to a stable complex-locked state, i.e., a stable fixed point in complex vector space defined in Ref. Thümler et al., 2023.
For generic complex coupling $\alpha=\frac{\pi}{2}-0.2$ (purple circles in Fig. 5) the system exhibits a discontinuous synchronization transition at substantially smaller $|K|$, where the majority of oscillators become closely locked and $r$ is close to one. Past the transition, pronounced synchrony emerges with a distribution variance in the real parts $x_{n}$ on the order of $10^{-3}$ and non-zero imaginary parts. Finally, for a purely imaginary coupling with $\alpha=\frac{\pi}{2}$, the oscillators are incoherently drifting, yielding a low order parameter relative to other scenarios even at substantially large $|K|$ (green open squares). These findings qualitatively align with those for the two-oscillator systems discussed in Sec. III.3, where the systems fail to exhibit any classical locked state and are asynchronous for purely imaginary coupling, whereas oscillators subject to a generic complex coupling become locked at a stable fixed point and exhibit a substantial degree of synchrony.
@@ V Conclusions and Open Questions
More specifically, we have presented a range of results: For $N=2$ coupled units and purely imaginary coupling, we discovered motions of libration around fixed points of in-phase and antiphase states which are respectively local minima and maxima of a conserved quantity, and a ceaseless rotation along the cylindrical axis with an infinite rotation number $q=\infty$. No attractive locked state emerges even for arbitrarily large $|K|$. In contrast, for generic complex coupling, trajectories rotate a few times, exhibit finite rotation numbers $q<\infty$, and stop at a stabilized fixed point. We have analytically determined the rotation number and derived its asymptotic scaling as $\alpha\rightarrow\pi/2$ from below, where the coupling becomes purely imaginary. For small $\beta=\pi/2-\alpha$, synchrony stabilizes: Slightly turning on the real part of a coupling, the neutrally stable fixed points of in-phase synchrony become stabilized to first order in $\beta$. Larger, finite-size systems of $N>2$ units similarly exhibit a classically locked state for purely imaginary coupling whereas a discontinuous transition emerges to a locked state for a generic complex coupling. The analyses above catalyze a range of future research.
@@
Acknowledgements.
Partially supported by German Federal Ministry for Education and Research (BMBF) under grant number 03SF0769 (ResiNet) as well as the German National Science Foundation (Deutsche Forschungsgemeinschaft DFG) under grant number DFG/TI 629/13-1 and under Germany´s Excellence Strategy – EXC-2068 – 390729961, through the Cluster of Excellence Physics of Life.